Marginal Analysis: Suppose factor prices for labour and capital are w = 4 and r = 8. At the observed amount of capital and labour the marginal product of capital is observed to be 15 and the marginal product of labour 5. Is the firm profit maximizing? Explain how it could increase its profits.
Answer #1

Firm is maximizing profits if Marginal product of labor/w = Marginal product of capital/r
Marginal product of labor/w = 5/4 = 1.25
Marginal product of capital/r = 15/8 = 1.875
So, the firm is not maximizing profits as Marginal product of labor/w < Marginal product of capital/r. So, to increase its profits, firm should increase its use of capital and decrease its use of labor.